# Resurgent structure of 2d Yang-Mills theory on a torus

### v2

@@ 1 Introduction
However, it is not known how to formulate the large $N$ duality precisely when $N$ is finite. At large but finite $N$, the factorization is only schematic and it needs to be modified by summing over RR fluxes over the finite $T^{2}$ Vafa:2004qa ; Dijkgraaf:2005bp . More importantly, while on the one side the partition function of 2d Yang-Mills is non-perturbatively well-defined for any finite $N$, on the other side the topological string partition function is only defined perturbatively. One therefore has to find its non-perturbative completion, taking into account appropriate non-perturbative corrections of the order $\mathcal{O}({\rm e}^{-N})$. There have been various attempts to find these non-perturbative corrections Lelli:2002gr ; Matsuo:2004nn ; deMelloKoch:2005rq ; Dhar:2006ru ; Okuyama:2018clk . The most promising proposal is given by Okuyama and Sakai Okuyama:2018clk , based on free fermion formulation of the 2d Yang-Mills Minahan:1993np ; Douglas:1993wy . However, only the 1-instanton amplitude of the proposal has been tested, and its multi-instanton amplitudes are not only un-tested, but in fact suffer from some limitations.
@@
In this paper, we apply the same method to find non-perturbative corrections to the partition function of the topological string dual to 2d Yang-Mills. We find closed form formulas for both single and multi-instanton amplitudes up to arbitrary instanton orders, and they turn out to be slightly different from the formulas in the usual topological string Gu:2022sqc ; Gu:2023mgf . Our 1-instanton amplitude agrees with that of Okuyama and Sakai Okuyama:2018clk , but our multi-instanton amplitudes are different. They pass high precision numerical tests from resurgence theory, and they overcome the limitations suffered by previous proposals. In particular, our formulas are valid for non-vanishing $\theta$ as well. Based on these results, we propose the non-perturbative partition function,
$\mathcal{Z}^{\text{np}}(t;g_{s};\bm{\sigma})=\sum_{n=0}^{\infty}\frac{1}{n!}B_{n}\left(\left\{\sigma_{j}\pm\frac{{\mathsf{i}}}{2}j!\sqrt{\frac{2\pi}{g_{s}}}\frac{(-1)^{j}}{j^{3/2}}\right\}_{j=1,2,\ldots,n}\right)\mathscr{S}^{(\pm)}\mathcal{Z}(t+ng_{s};g_{s}),$ (1.1)
taking into account contributions from all the real instantons, and it is real when both the modulus $t$ and the string coupling $g_{s}$ are positive. It involves an infinite family of real parameters $\bm{\sigma}=(\sigma_{1},\sigma_{2},\ldots)$, and in the simplest case $\bm{\sigma}=\bm{0}$, it is just the medium resummation of the perturbative partition function. Here $B_{n}(\ldots)$ are complete Bell polynomials, and $\mathscr{S}^{(\pm)}$ denote the lateral Borel resummations. Note that one can choose either of the two lateral Borel resummations $\mathscr{S}^{(\pm)}$ (see (4.10) for definition) with the same sign $\pm$ in the Bell polynomials to match it. We believe this should be used to formulate the precise duality between 2d Yang-Mills on torus and topological string for finite $N$.
In addition, we also explore complex instantons, and find two infinite towers of complex instanton sectors. They should correspond to non-trivial BPS D-brane bound states in type II string. We also discuss a little the wall-crossing behavior of these instanton sectors.
@@ 2.1 String interpretation
Here $\operatorname{tr}$ is the trace in the fundamental representation. The partition function can be calculated by summing over representations $R$ of the gauge group Migdal:1975zg ; Rusakov:1990rs
@@ 2.2 Perturbative free energies
To compute the free energy $\mathcal{F}_{g}(t)$ efficiently, we begin with a high-order, e.g., 2000th order, $q$-series expansion of $e_{\ell}$. Utilizing the recursive relation
@@
derived from (2.14), the $q$-series expression of $h_{n}^{\text{top}}$ can be computed efficiently. Substituting $h_{n}^{\text{top}}$ into the recursive relation (2.13), we are able to obtain the 2000th-order $q$-series expression of $\mathcal{Z}_{n}^{\text{top}}(t)$ for $n\leq 200$. Since $\mathcal{Z}_{n}^{\text{top}}(t)$ is finitely generated by Eisenstein series, a sufficiently high-order $q$-series expansion allows one, in principle, to recover the modular expression of $\mathcal{Z}_{n}^{\text{top}}(t)$.
@@
where the holomorphic part $P_{0}(z)$ is the holomorphic free energy $\mathcal{F}_{g}(t)$.
@@ 2.3 Non-perturbative proposal
It was argued in Vafa:2004qa that the factorization (2.4) is only schematic and to make it exact one has to sum over RR-fluxes over the torus, arriving at
@@
An even more refined version is proposed in Dijkgraaf:2005bp . It is, however, difficult to generalize formulas like (2.29) to cases with large but finite $N$. As argued in Okuyama:2018clk , in this case the meaning of (2.29) is unclear: the left hand side is well-defined for finite $N$ and thus finite $g_{s}$, while the right hand side is only defined as asymptotic perturbative power series of $g_{s}$.
An attempt to solve this problem was made in Okuyama:2019rqn , in which process, a non-perturbative version of topological string free energy was proposed. The partition function of the 2d Yang-Mills theory allows a free fermion formulation Minahan:1993np ; Douglas:1993wy
@@
where $p={\rm e}^{-g_{s}}$. Let us focus on the simple scenario $\theta=0$ and assume that $N$ is always even. The partition function reduces to
@@
More explicitly, one can show that Okuyama:2018clk
@@
However, there are several problems with this proposal. First of all, even though this proposal includes corrections from multi-instanton amplitudes to arbitrary high orders, only the 1-instanton amplitude was tested Okuyama:2018clk . Second, this proposal was derived with the assumption that $N$ is even and $\theta=0$, which is simply too restrictive. On the one hand, when $N$ is odd, (2.30) seems to indicate that in the generating series (2.37) $\vartheta_{2}$ should be replaced by $\vartheta_{3}(p)$, which nonetheless cannot be distinguished by the 1-instanton test, as in the $g_{s}\rightarrow 0$ limit,
@@ 4.5 Non-perturbative proposal
This is also verified numerically in Tabs. 4.2, 4.3, 4.4, 4.5: as higher order instanton corrections are included, both of the two lateral resummations have progressively smaller imaginary parts, and they converge to the same real number. Note that the simplest member of this family is the celebrated medium resummation of the perturbative partition function

### v3

@@ 1 Introduction
However, it is not known how to formulate the large $N$ duality precisely when $N$ is finite. At large but finite $N$, the factorization is only schematic and it needs to be modified by summing over RR fluxes over the finite $T^{2}$ Vafa:2004qa ; Dijkgraaf:2005bp . More importantly, while on the one side the partition function of 2d Yang-Mills is non-perturbatively well-defined for any finite $N$, on the other side the topological string partition function is only defined perturbatively in the large $N$ expansion. One therefore has to find its non-perturbative completion, taking into account appropriate non-perturbative corrections of the order $\mathcal{O}({\rm e}^{-N})$. These are spacetime instantons on the string theory side due to D-branes, and they correspond to large $N$ instanton corrections on the 2d YM theory side. There have been various attempts to find these non-perturbative corrections Lelli:2002gr ; Matsuo:2004nn ; deMelloKoch:2005rq ; Dhar:2006ru ; Okuyama:2018clk . The most promising proposal is given by Okuyama and Sakai Okuyama:2018clk , based on free fermion formulation of the 2d Yang-Mills Minahan:1993np ; Douglas:1993wy . However, only the 1-instanton amplitude of the proposal has been tested, and its multi-instanton amplitudes are not only un-tested, but in fact run into various problems of inconsistency, especially for non-vanishing $\theta$ angles.
@@
In this paper, we apply the same method to find non-perturbative corrections to the partition function of the topological string dual to 2d Yang-Mills. We find closed form formulas for both single and multi-instanton amplitudes up to arbitrary instanton orders, and they turn out to be slightly different from the formulas in the usual topological string Gu:2022sqc ; Gu:2023mgf . Our 1-instanton amplitude agrees with that of Okuyama and Sakai Okuyama:2018clk , but our multi-instanton amplitudes are different. They pass high precision numerical tests from resurgence theory, and they overcome the limitations suffered by previous proposals. In particular, our formulas are valid for non-vanishing $\theta$ as well. Based on these results, we propose the non-perturbative partition function (4.93) which take into account contributions from all real instantons. Even though it does not seem to be unique on the surface and depends on an infinite family of real parameters, we further conjecture that there is a canonical choice of these parameters i.e., (4.96), which should be used to formulate the precise duality between 2d Yang-Mills on torus and topological string for large but finite $N$.
As we mentioned, the non-perturbative corrections that we have consistently included are large $N$ instanton corrections for 2d YM theory, or worldsheet instanton corrections for the dual topological string theory due to D-branes, and these are all real instantons with real instanton actions. There could equally be complex instantons with complex instanton actions. Even though they do not play an important role in the non-perturbative partition function of the chiral 2d YM theory, they do correspond to important generic D-branes in the dual topological string theory with complex central charges. In this paper, we also explore complex instantons, and we find two infinite towers of complex instanton sectors. They should correspond to non-trivial BPS D-brane bound states in type II string. We also discuss the wall-crossing behavior of these instanton sectors.
@@ 2.1 String interpretation
Here $\operatorname{tr}$ is the trace in the fundamental representation. The partition function can be calculated by summing over representations $R$ of the gauge group Migdal:1975zg ; Rusakov:1990rs ; Witten:1991we
@@ 2.2 Perturbative free energies
To compute the free energy $\mathcal{F}_{g}(t)$ efficiently, we begin with a high-order, e.g., 1000th order, $q$-series expansion of $e_{\ell}$. Utilizing the recursive relation
@@
derived from (2.14), the $q$-series expression of $h_{n}^{\text{top}}$ can be computed efficiently. Substituting $h_{n}^{\text{top}}$ into the recursive relation (2.13), we are able to obtain the 1000th-order $q$-series expression of $\mathcal{Z}_{n}^{\text{top}}(t)$ for $n\leq 200$.  [^6] The data and the program to generate these data are put on the website https://github.com/cloud-atlas-47/2dYM-np. Since $\mathcal{Z}_{n}^{\text{top}}(t)$ is finitely generated by Eisenstein series, a sufficiently high-order $q$-series expansion allows one, in principle, to recover the modular expression of $\mathcal{Z}_{n}^{\text{top}}(t)$. For instance, we obtin the full modular expression for $\mathcal{Z}_{n}^{\text{top}}(t)$ up to $n=60$.
@@
where the holomorphic part $P_{0}(z)$ is the holomorphic free energy $\mathcal{F}_{g}(t)$. Using this method, we have calculated the non-holomorphic free energy $F_{g}$ up to $g=60$.  [^8] The data and the program to generate these data are put on the website https://github.com/cloud-atlas-47/2dYM-np.
@@ 2.3 Non-perturbative proposal
It was pointed out in Vafa:2004qa that the factorization (2.4) was only approximately valid for large but finite $N$, and one had to take into account non-perturbative corrections $\mathcal{O}({\rm e}^{-N})$ by summing over RR-fluxes over the torus, arriving at
@@
An even more refined version was proposed in Dijkgraaf:2005bp . It was later argued in Okuyama:2018clk that even these proposals were flawed: e.g., the left hand side of (2.29) was well-defined for finite $N$ and thus finite $g_{s}$, while the right hand side of (2.29) was only defined as asymptotic perturbative power series of $g_{s}$.
An attempt to solve this problem was made in Okuyama:2018clk , in which process, a non-perturbative version of topological string free energy was proposed. The partition function of the 2d Yang-Mills theory allows a free fermion formulation Minahan:1993np ; Douglas:1993wy
@@
where $p={\rm e}^{-g_{s}}$. Restricted to the simple scenario of $\theta=0$ and assuming that $N$ is even, the partition function reduces to
@@
More explicitly, it was shown in Okuyama:2018clk that
@@
Let us comment that there are several problems with this proposal. First of all, even though this proposal includes corrections from multi-instanton amplitudes to arbitrary high orders, only the 1-instanton amplitude was tested Okuyama:2018clk . Second, this proposal was derived with the assumption that $N$ is even and $\theta=0$, which is simply too restrictive. On the one hand, when $N$ is odd, (2.30) seems to indicate that in the generating series (2.37) $\vartheta_{2}$ should be replaced by $\vartheta_{3}(p)$, which nonetheless cannot be distinguished by the 1-instanton test, as in the $g_{s}\rightarrow 0$ limit,
@@ 4.5 Non-perturbative proposal
This is also verified numerically in Tabs. 4.2, 4.3, 4.4, 4.5: as higher order instanton corrections are included, both of the two lateral resummations have progressively smaller imaginary parts, and they converge to the same real number.
We note that the formula for non-perturbative partition function (4.93) is not unique as it depends on an infinite family of free parameters $\bm{\sigma}$. It thus seems that the resurgence analysis does not capure the complete information of non-perturbative corrections. We make two comments here. Firstly, it is not uncommon that the resurgence analysis fails to uncover the full non-perturbative corrections. For instance, in quantum mechanical models whose potentials are symmetric, resurgence analysis does not capture non-perturbative information in the energy spectrum that arise from (quantum mechanically breaking) these symmetries, such as the removal of energy degeneracy in the double-well model with parity symmetry, and the Bloch angle in the cosine model with periodic symmetry (see for instance discussion in Marino2015:instantons ). Second, nevertheless, there does not seem to be any additional symmetries in 2d YM which are quantum mechanically broken, parameterized by some order parameter. Therefore, we conjecture that we can simply take the simplest member in the family (4.93), which is the celebrated medium resummation of the perturbative partition function